OpenStudy (anonymous):

. Find the product: (8x-3y)²

OpenStudy (anonymous):

(8x-3y)(8x-3y) Know how to FOIL?

OpenStudy (anonymous):

(8x-3y)(8x-3y) 64x²-24xy+24xy-9y² 64x²-48xy-9y²

OpenStudy (anonymous):

Close. A negative times a negative is positive, so (-3y)(-3y) = +9y^2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Also, you can use the Binomial Theorem to know that in general, (a+b)^2 = a^2+2ab+b^2 So (ax+by)^2 = a^2x^2+2abxy+b^2y^2 And (ax-by)^2 = a^2x^2-2abxy+b^2y^2